# Boundary Exploration of Next Best View Policy in 3D Robotic Scanning

### v1

@@ IV-E Performance Comparison
We compare our proposed method with approaches that utilize point clouds as input data, including PC-NBV [16], a deep network for next-best-view planning, and SEE [12], a density-based edge exploration method for NBV planning. For the comparison, we use the pretrained model provided by PC-NBV and the default parameter settings for small objects as outlined in SEE’s work. Additionally, the following policies are used as baselines:
@@
    Random Boundary: The next best view is randomly chosen from the detected boundary of the retrieved point cloud.
@@
  3. 3.
    Random Uniform Sphere (Random U- Sphere): The next camera position is randomly selected from uniformly distributed points on a sphere.
  4. 4.
    Ours (BENBV): It searches and selects the optimal next view given the detected boundary points..
@@
    Ours via Deep Learning (BENBV-Net): Using the detected boundary and the captured point cloud, it predicts the score of the next best view directly without a search process.
As shown in Tables I, II, and III, we evaluate the performance of our NBV policy in comparison to other methods on the ShapeNet, ModelNet, and 3D Repository datasets. Specifically, each method is run 60 times on the 3D Repository dataset and twice on ShapeNet and ModelNet. The initial view is provided randomly on a spherical surface. The reported coverage represents the final coverage achieved after 15 scanning attempts, while the overlap reflects the overlap obtained during the first five scans. The best results are marked in bold and second best result are underlined.
The results show that our method achieves final coverage rates of 89% and 95% on the test datasets, surpassing all compared approaches. The scanning efficiency is approximately 9.0 and 13.0, outperforming other traditional methods. Additionally, our approach maintains an average overlap of 50% to 60%, significantly higher than the 30% achieved by PC-NBV. The overlap and efficiency of the random boundary method are higher than those of other random-based methods, indicating that exploring based on the boundary better considers overlap and final coverage compared to random methods that limit the view space. Finally, the Chamfer and Hausdorff distances used to evaluate the quality of the reconstructed 3D scans are either lower than or comparable to those of other methods, further validating the accuracy of our approach.
TABLE I: Evaluation results of Next-Best-View policies on ShapeNet, the number of views is set to 15 maximum. The symbols $\uparrow$ and $\downarrow$ indicate whether a higher or lower value is better, respectively.
@@
BENBV 89.09 59.29 0.38 9.25 8.76
Random Boundary 86.27 54.38 0.39 10.44 6.58
Random Sphere 86.01 36.03 0.36 10.06 6.34
Random U- Sphere 86.41 34.62 0.36 10.05 6.48
PC-NBV 87.36 33.81 0.34 9.78 7.18
SEE 62.89 55.21 1.26 17.62 4.13
BENBV-Net 85.92 55.25 0.51 10.91 7.51
@@
BENBV 89.07 62.21 0.39 8.81 9.01
Random Boundary 87.77 53.80 0.32 8.86 7.12
Random Sphere 87.87 37.04 0.28 8.30 6.97
Random U- Sphere 87.31 34.30 0.29 8.53 6.81
PC-NBV 88.19 33.05 0.27 8.21 7.49
SEE 65.80 56.22 1.24 17.29 4.40
BENBV-Net 87.34 58.17 0.41 9.43 8.03
@@
BENBV 95.04 53.52 0.11 5.18 13.00
Random Boundary 93.82 57.14 0.12 6.13 9.03
Random Sphere 87.88 32.31 0.21 7.11 6.52
Random U- Sphere 85.49 24.97 0.25 7.45 5.82
PC-NBV 91.93 32.57 0.14 6.22 8.59
SEE 77.73 57.87 0.50 12.54 5.53
BENBV-Net 94.25 47.01 0.12 5.90 11.23
Using 50%, 80%, and 90% coverage as milestones for evaluating the NBV policy, the number of views required to reach each milestone is detailed in Table IV. The results indicate that random-based methods, such as Random Uniform Sphere and PC-NBV, achieve 50% coverage within four scans but require significantly more scans, such as at least 12 for the ShapeNet dataset and 10 for the 3D Repository, to reach 90% coverage. In contrast, BENBV method achieves these milestones more efficiently. For instance, it requires approximately 6 views to reach 80% coverage and 7 views to achieve 90% coverage on dataset 3D repository. Additionally, BENBV-Net demonstrates comparable performance, achieving higher convergence than PC-NBV with fewer scans. Therefore, the results validate the effectiveness of our method, including the model-based approach and BENBV-Net, in rapidly achieving high coverage with fewer scanning views.
@@
ShapeNet ModelNet40 3D Repository
NBV Policy Coverage
@@
BENBV 4.70 8.21 10.17 4.41 8.10 9.89 3.83 6.40 7.31
Random Boundary 4.64 10.66 13.11 4.27 9.89 12.32 3.76 8.31 10.39
Random Sphere 4.59 10.90 13.57 4.33 10.01 12.60 4.69 10.59 13.48
Random U- Sphere 4.52 10.60 13.34 4.40 10.10 12.81 5.34 11.85 14.69
PC-NBV 3.55 9.12 12.17 3.53 8.90 11.78 3.23 7.61 10.70
SEE 7.23 14.03 15.22 6.34 13.12 14.94 4.03 10.46 14.05
BENBV-Net 4.38 8.91 11.44 4.12 8.52 10.88 3.20 6.50 8.39
Furthermore, the detailed overlap metrics are outlined in Table V, where we present the overlap during the first 3 scans, scans 3 to 6, and scans 6 to the final attempt as intervals. According to the results, BENBV generally achieves the highest overlap performance, such as 54% on ShapeNet and 57% on ModelNet for the first 3 scans. In addition, BENBV-Net achieves the second-best overlap, with 49% and 52% on the same datasets. The PC-NBV method has the lowest overlap during the first 3 scans, with averages of 24%, 48%, and 66% on the tested datasets. However, BENBV-Net can achieve an average overlap of 46%, 65%, and 80% on the given datasets. It demonstrates that the effectiveness of PC-NBV lies in its approach of ignoring overlap and focusing solely on the high unseen regions during the view selection process. The SEE method achieves the highest overlap between scans 3 and 6 but struggles to identify views that significantly enhance coverage. This is evident in Table IV, where SEE exhibits the lowest scanning efficiency.
TABLE V: Comparison of NBV policies based on overlap retrieved during the first 3 scans, scans 3 to 6, and scans 6 to the final view.
ShapeNet ModelNet40 3D Repository
NBV Policy Scan View
1 - 3 3 - 6 6 - 15 1 - 3 3 - 6 6 - 15 1 - 3 3 - 6 6 - 15
BENBV 54.95 68.38 85.48 57.40 71.59 85.64 48.45 65.85 86.11
Random Boundary 49.21 63.29 72.37 48.79 63.33 71.64 51.50 67.06 77.75
Random Sphere 28.91 49.33 61.37 30.80 47.68 59.50 25.01 45.43 59.06
Random U- Sphere 27.50 47.36 58.53 27.82 45.37 57.56 18.80 36.22 51.43
PC-NBV 24.49 49.04 65.41 25.90 45.74 63.81 22.68 50.50 69.22
SEE 44.69 72.33 76.20 45.28 73.94 79.13 44.90 78.36 82.55
BENBV-Net 49.62 66.48 81.53 52.47 68.67 81.54 39.09 62.73 79.38
To provide a clearer understanding of the compared NBV policies, the convergence of each attempt and the overlap between views are illustrated using the 3D repository dataset as an example, as shown in Figure 5 and Figure 6. These figures demonstrate that our method effectively considers both overlap and coverage. First, BENBV achieves the highest coverage compared to other methods, except for the Blade data, which has an internal composition that limits observability and prevents further coverage improvement. Second, both the model-based method and BENBV-Net consider overlap. BENBV-Net performs better in terms of overlap for the Statue of Liberty dataset, demonstrating its ability to generalize to novel objects, which is a challenge even for the model-based method.
Figure 5: Coverage for each view in the 3D Repository, with the first view (0-th view) initialized by random scanning based on a sphere surface.
@@
Figure 6: Overlap during the first 3 scans, scans 3 to 6, and scans 6 to final in the 3D Repository dataset, where the methods are labeled as follows: 0: BENBV-Net; 1: SEE; 2: PC-NBV; 3: Random uniform sphere; 4: Random sphere; 5: Random boundary; 6: BENBV.
@@
A significant drawback of the model-based method is its increased processing time, as illustrated in Table VI, which presents the average time spent scanning each object in the tested datasets. The results reveal that the model-based approach (BENBV) requires approximately 8 to 13 times more time than other methods. In contrast, our learning-based approach, BENBV-Net, takes only 7.8 seconds, demonstrating comparable efficiency in terms of time consumption.
@@
TABLE VI: The average process time (seconds) for each object in tested datasets.
Method
@@
In summary, compared to the tested random-based approaches, SEE, and PC-NBV, the random-based methods can achieve moderate final coverage but require more views to reach 90% coverage. PC-NBV achieves high coverage rapidly but neglects coverage during data acquisition and shows the lowest overlap among the tested methods. SEE provides high overlap during scanning but struggles to identify views that significantly increase coverage. The experiments demonstrate that our method achieves higher completeness more quickly while also considering the overlap between views.
A more specific demonstration is illustrated in Figure. 7, where the object shown is the bunny. The same initial view is provided, with the initial coverage at 28%. The coverage (C) and overlap (O) are shown at the 3rd, 6th, and 10th scans. It demonstrates that BENBV-Net achieves the highest overlap, while maintaining a coverage higher than the other methods.
Figure 7: An example of the NBV policies is illustrated using a bunny object.
@@ V Conclusion and Future Work
In this paper, we propose an NBV policy that explores the boundary of the point cloud while balancing coverage and overlap. We present two approaches: a model-based approach, BENBV, which uses a reference model to calculate scores for potential views based on overlap and coverage, selecting the view with the highest score as the NBV. The second approach, BENBV-Net, is learning-based and uses a deep learning network trained on previously collected data. It predicts scores directly from the scanned data and proposed views, allowing NBV selection without the need for a reference model.
The experiments demonstrate that both of our proposed methods achieve high coverage and overlap when evaluated on public datasets such as ShapeNet, ModelNet40, and the 3D Repository. Furthermore, the time efficiency of our learning-based method is comparable to traditional approaches, underscoring its practicality for 3D scanning tasks.

### v2

@@ IV-E Performance Comparison
We compare our proposed method with several approaches that utilize point clouds as input data, including a learning-based NBV prediction method and a density-based edge exploration method. Specifically, the evaluated methods are:
@@
    PC-NBV [16]: A deep neural network for next-best-view planning. The pretrained model provided by the authors is used in this work.
@@
    SEE [12]: A surface-edge exploration method that selects views based on point cloud density and edge characteristics. The default parameter settings for small objects are applied in this work.
  3. 3.
    Random Boundary: The next-best view is randomly selected from the detected boundary of point cloud.
  4. 4.
@@
    Random Uniform Sphere (Random U-Sphere): The next camera position is randomly selected from a set of uniformly distributed points on a sphere.
  6. 6.
    Ours (BENBV): Our method searches for and selects the optimal next view given the detected boundaries.
  7. 7.
    Ours via Deep Learning (BENBV-Net): Given the detected boundary and captured point cloud, this model directly predicts the next-best view score without performing an explicit search.
Each method is executed 60 times on the 3D Repository dataset and twice on the ShapeNet and ModelNet datasets. For all tested methods, the initial view is randomly selected from a spherical surface. The performance of our NBV policy in comparison with other methods on the ShapeNet, ModelNet, and 3D Repository datasets is summarized in Tables I, II, and III, respectively. The reported coverage denotes the final coverage achieved after 15 scans. The best results are highlighted in bold, and the second-best results are underlined.
The results show that our method (BENBV) achieves the highest final coverage across all compared approaches. More importantly, it reaches scanning efficiency scores of over 10.0 and 13.0, outperforming all other methods. In addition, our approach consistently selects views with stable overlap, reaching a value of 8.16 on the 3D Repository dataset. The Chamfer and Hausdorff distances, used to evaluate reconstruction quality, are either lower than or on par with those of the other methods, supporting the accuracy of our approach. It is also worth noting that the random boundary method achieves a comparable CV of overlap, suggesting that selecting views along object boundaries is a promising direction for NBV policy design. However, its low efficiency remains a significant drawback.
Regarding BENBV-Net, it generally delivers the best performance among the model-free methods, including PC-NBV and SEE. On the ModelNet40 dataset, PC-NBV achieves lower Chamfer and Hausdorff distances than our method, which is commendable. However, its CV of overlap is the highest (i.e., most unstable), and its scanning efficiency is also lower than that of BENBV-Net. Similarly, on the ShapeNet dataset, PC-NBV again achieves a lower Chamfer distance, but both its efficiency and overlap stability are worse compared to our method. As for SEE, which is based on edge density, it shows comparable overlap stability only on the 3D Repository dataset. On ModelNet and ShapeNet, however, it performs the worst among all evaluated methods. As a result, our proposed methods, BENBV and BENBV-Net, achieve the best and second-best performance in terms of efficiency, final coverage, and overlap stability. Although the final reconstruction quality does not surpass that of PC-NBV, it remains on a comparable level.
TABLE I: Evaluation results of Next-Best-View policies on ShapeNet.
@@
BENBV 91.27 12.58 0.27 8.62 9.22
Random Boundary 86.64 34.30 0.38 10.34 6.68
Random Sphere 86.10 43.80 0.36 9.94 6.42
Random U- Sphere 86.80 48.95 0.35 10.05 6.54
PC-NBV 87.36 50.80 0.33 9.72 7.11
SEE 63.84 41.46 1.24 18.11 4.20
BENBV-Net 88.24 23.19 0.35 9.62 7.71
@@
BENBV 92.58 15.33 0.22 7.01 10.16
Random Boundary 87.78 33.42 0.32 9.18 7.10
Random Sphere 88.3 47.90 0.27 8.38 7.08
Random U- Sphere 87.93 49.91 0.28 8.32 6.92
PC-NBV 88.24 52.41 0.27 8.23 7.44
SEE 67.82 38.86 1.06 16.52 4.61
BENBV-Net 89.18 26.18 0.31 8.45 8.34
@@
BENBV 95.70 8.16 0.09 5.20 13.45
Random Boundary 94.28 23.55 0.11 5.91 9.5
Random Sphere 91.24 33.27 0.15 6.56 7.55
Random U- Sphere 91.97 34.45 0.14 6.47 7.74
PC-NBV 92.11 34.76 0.14 6.19 8.76
SEE 77.67 27.33 0.52 12.34 5.76
BENBV-Net 94.44 18.42 0.11 5.65 10.64
To better understand the achieved final coverage, we use the number of views required to reach 50%, 80%, and 90% coverage as key milestones for evaluating NBV policies. These results are detailed in Table IV. Random-based methods typically reach 50% coverage within four scans; however, they require significantly more views to reach higher coverage, around 13 scans for ShapeNet and 10 for the 3D Repository dataset to reach 90%. In contrast, our BENBV method achieves these milestones more efficiently. For example, it reaches 80% coverage with approximately 6 views and 90% with just 7 views on the 3D Repository dataset. BENBV-Net shows similarly strong performance, converging faster than PC-NBV while requiring fewer scans.
It is also worth noting that PC-NBV can reach 50% coverage with only 3 scans, but this comes at the cost of significantly lower overlap compared to our approach. Since overlap stability is crucial in practical 3D scanning tasks, this limits its overall usefulness. These results highlight a key distinction: while existing methods focus solely on maximizing final coverage, they often overlook the importance of overlap stability. Our method balances both goals: high coverage with fewer scans and stable overlap in the initial scanning phase.
@@
NBV Policy ShapeNet ModelNet40 3D Repository
@@
BENBV 3.94 7.67 9.90 3.76 7.02 9.11 3.26 6.30 7.11
Random Boundary 4.54 10.69 12.98 4.24 9.83 12.36 3.78 8.12 9.93
Random Sphere 4.55 10.81 13.41 4.13 9.82 12.48 3.92 9.13 12.09
Random U- Sphere 4.49 10.48 13.27 4.23 9.97 12.72 3.83 8.74 11.88
PC-NBV 3.53 9.14 12.29 3.47 8.87 11.86 3.22 7.49 10.52
SEE 6.66 13.81 15.21 6.28 12.84 14.71 4.17 10.71 13.47
BENBV-Net 4.30 8.93 11.44 4.11 8.34 10.70 3.42 6.94 8.87
Furthermore, the detailed overlap metrics are presented in Table V, showing the average overlap during the first three scans. In the table, the lowest overlap values, often considered unexpected, are marked with a box for emphasis. According to the results, BENBV consistently achieves stable overlap performance, maintaining around 43% overlap across the first three scans. BENBV-Net also demonstrates similarly stable overlap, averaging above 40%.
In contrast, PC-NBV shows some of the lowest overlap values, with averages of 11.92%, 23.72%, and 38.13%, respectively. This helps explain why PC-NBV is able to reach 50% coverage within just 3 scans (as shown in Table IV): it prioritizes coverage while largely sacrificing overlap. In addition, while SEE achieves higher overlap than PC-NBV, Random Sphere, and Random U-Sphere, it still only reaches about half the overlap achieved by our methods, BENBV and BENBV-Net.
Interestingly, the random boundary method consistently achieves relatively high overlap in the first scan. This suggests that exploring boundaries may be an effective strategy for achieving reliable overlap in NBV planning.
TABLE V: Comparison of NBV policies based on the first 3 scans.
NBV Policy ShapeNet ModelNet40 3D Repository
1 th 2 nd 3 rh 1 th 2 nd 3 rh 1 th 2 nd 3 rh
BENBV 43.05 43.23 44.68 43.42 43.90 46.30 43.33 42.65 41.88
Random Boundary 39.42 51.71 57.33 41.37 50.33 56.66 42.01 52.55 59.04
Random Sphere 19.04 29.73 39.85 19.18 32.54 39.59 19.11 30.78 40.95
Random U- Sphere 17.61 29.05 36.84 18.92 29.45 36.27 19.22 29.55 39.85
PC-NBV 11.92 23.72 38.13 16.33 24.70 38.21 10.07 22.21 35.34
SEE 20.21 51.64 67.15 24.00 53.99 67.70 20.23 46.45 65.59
BENBV-Net 40.45 43.13 43.40 40.15 43.19 46.55 43.82 43.18 45.04
To provide a clearer comparison of the NBV policies, we illustrate the convergence behavior and overlap between views using the 3D Repository dataset as an example, as shown in Figure. 5 and Figure. 6. As seen in Figure. 5, BENBV consistently achieves the highest coverage across most cases. BENBV-Net performs similarly, with the exception of the Statue of Liberty sample, where it falls slightly behind. Based on Figure 6, both the model-based method and BENBV-Net exhibit stable overlap performance. In some cases, such as the Bunny and Horse datasets, BENBV-Net performs on par with BENBV. Moreover, since the 3D Repository dataset was not used during training, these results also highlight BENBV-Net’s ability to generalize to unseen objects which is a challenging task even for model-based methods.
Figure 5: Coverage for each view for the 3D Repository dataset.
@@
Figure 6: CV of overlap over all scans in the 3D Repository Dataset. Methods are labeled as follows: 0: BENBV, 1: Random Boundary, 2: Random Sphere, 3: Random Uniform Sphere, 4: PC-NBV, 5: SEE, 6: BENBV-Net.
@@
A significant drawback of BENBV is its longer processing time, as shown in Table VI, which reports the average time required to scan each object across the tested datasets. The results indicate that BENBV takes approximately 8 to 13 times longer than other model-free methods. In contrast, our learning-based approach, BENBV-Net, requires only 7.8 seconds per object, demonstrating competitive efficiency in terms of runtime.
@@
TABLE VI: The average process time (s) for each object in tested datasets.
@@
In summary, the random-based approaches achieve moderate final coverage but require more views to reach 90% completeness. PC-NBV reaches high coverage quickly but largely overlooks overlap during data acquisition. SEE, on the other hand, ensures high overlap during scanning but struggles to select views that meaningfully increase coverage. Our experiments show that the proposed method achieves higher coverage more efficiently while maintaining stable overlap between views, particularly during the early stages of scanning. In other words, it strikes a balanced trade-off between overlap acquisition and final coverage.
@@ V Conclusion and Future Work
In this paper, we propose an NBV policy that explores the boundary of the point cloud while balancing coverage and overlap. We present two approaches: a model-based method, BENBV, which uses a reference model to search for potential views based on a carefully designed strategy, selecting the view with the highest score as the NBV; and a learning-based method, BENBV-Net, which employs a deep neural network trained on data generated by the model-based approach. BENBV-Net predicts scores directly from the accumulated data, allowing NBV selection without reference models.
The experiments demonstrate that both of our proposed methods achieve high coverage and stable overlap when evaluated on public datasets, including ShapeNet, ModelNet40, and the 3D Repository. Most importantly, the scanning efficiency ranks as the best or second-best among all tested methods. Furthermore, the time efficiency of our learning-based approach is comparable to traditional methods, highlighting its practicality for real-world 3D scanning tasks.